OpenStudy (bloomlocke367):

Is it possible to have a piecewise function with only part of a circle?.. I know the equation isn't a function so I'm not sure if it's an actual thing...

OpenStudy (anonymous):

I believe it's possible if you specify the domain.

OpenStudy (anonymous):

The only issue is the co-variant nature of that circle.

OpenStudy (bloomlocke367):

I tried and it didn't do it :O that's why I was wondering.. I have a project where I have to make a picture using piecewise functions and I only need part of a circle but I can't figure out how to do it.

OpenStudy (bloomlocke367):

@Nnesha

OpenStudy (bloomlocke367):

@Shalante

OpenStudy (anonymous):

I am too trying to figure out how piecewise function works for my brother

OpenStudy (anonymous):

It appears pretty intuitive until you figure out each specified domain must also account for different equations.

OpenStudy (bloomlocke367):

okay.. hmm

OpenStudy (bloomlocke367):

@King.Void.

OpenStudy (king.void.):

Not good at math

OpenStudy (bloomlocke367):

dang ok

OpenStudy (beginnersmind):

Think about the function \[f(x)= \sqrt{1-x^2}\] compare it with the equation of a circle of radius 1 centered on the origin: \[x^2+y^2 = 1\]

OpenStudy (bloomlocke367):

@Shalante

OpenStudy (jdoe0001):

you can have a function alone, and be only a semi-circle a circle is compounded of two semicircles really each semi-circle occurs when you use the square root the square root has a + and a - part, thus \(\large \pm \sqrt{\quad}\)

OpenStudy (bloomlocke367):

that only shows half of the circle..

OpenStudy (bloomlocke367):

yea okay XD I was about to say..

OpenStudy (bloomlocke367):

but I'm working on a piecewise project and I have to create a picture on a graph using piecewise functions and that's why I was wondering XD